# Quantized distributed Nash equilibrium seeking under DoS attacks

### v2

@@ V Simulation
For DoS attacks, we consider a sustained DoS attack with variable period and duty cycle, generated randomly. Over a simulation horizon of $1500$s, the DoS signal yields $|\Xi(0,1500)|=1465$s and $n(0,1500)=50$. This corresponds to values (averaged over $1500$s) of $\tau_{D}\approx 30$ and $1/T\approx 0.9767$, and hence $\Delta/\tau_{D}+1/T\approx 0.98$, which approaches the maximum tolerable DoS attack in Theorem 1. In Fig. 1, most of the simulation time is occupied by gray stripes implying the presence of DoS. The white stripes represent DoS-free intervals. Under such a strong level of DoS attacks, one can see in the top plot of Fig. 1 that the actions of the players converge to the NE. In the bottom plot in Fig. 1, one can see that the scaling function $\theta$ converges to zero as time elapses. Moreover, $\theta$ holds its value in the presence of DoS and decreases during DoS-free intervals, i.e., the zooming-in and holding. In Fig. 2, we present the quantized values for $x$ and $y$, i.e., (14) and (13), respectively. One can see that the actual utilized quantization range of $x$ is between $-10$ and $1$, and the one for $y$ is between $-6$ and $6$. They do not encounter saturation problem under DoS attacks and both satisfy the conditions for quantizer unsaturation in (39). One can see that the result in (39b) is conservative. One of the major reasons is the multiplication of $\|G\|$, $\|A_{0}\|$ and $\|S\|$ in (31b) for calculating the value in (39b). Note that constructiveness of quantization levels often occurs in quantized control of multi-agent systems even without DoS attacks (see e.g., [28, 29]). This is majorly due to the nature of distributed systems. If global state is unknown, the agents cannot “well” estimate the states of their neighbors. By contrast, for centralized systems, the problem of constructiveness can be well tackled [24, 25].
@@ VI Conclusions and future research
In the future, it is interesting to investigate quantized distributed NE seeking under DoS attacks by Lyapunov approaches by following [12]. Extending our results to the players in general linear form is also a practical direction[27]. It is also useful to apply our results to defend cyber-attacks [2]. At last, we expect that the results in our paper still hold under a directed and strongly connected graph in view of [6].
@@
$\leq\sqrt{1-2\delta\mu+(\delta)^{2}l^{2}}\|x(k)-x^{*}\|+N\delta l\|\bar{y}(k)\|$ (43)
@@
Therefore, (-B) describes the system dynamics between consecutive successful transmissions at $s_{r}$ and $s_{r+1}$, during which packet losses can occur due to DoS attacks.
@@
In (-B), $\|\pi(s_{0})\|$ is upper bounded by $\sqrt{N^{2}C_{x_{0}}^{2}+N(C_{x^{*}}+C_{x_{0}})^{2}}/\theta_{0}$. It is sufficient to calculate the upper bounds of $\|\beta(s_{0})\|$ and $\|\chi(s_{0})\|$ individually. We present the analysis by the following two cases. If $s_{0}=0$ (i.e., there has no DoS attack at the beginning), we have $\|\beta(s_{0})\|=\|\bar{y}(s_{0})\|/\theta(s_{0})=\|\bar{y}(0)\|/\theta(0)=\|y(0)-\mathbf{1}_{N}\otimes x(0)\|/\theta(0)=\|\mathbf{1}_{N}\otimes x(0)\|/\theta(0)\leq N\|\mathbf{1}_{N}\otimes x(0)\|_{\infty}/\theta(0)=NC_{x_{0}}/\theta(0)$. Similarly, one has $\|\chi(s_{0})\|=\|x(s_{0})-x^{*}\|/\theta(s_{0})\leq(\|x(s_{0})\|+\|x^{*}\|)/\theta(0)\leq\sqrt{N}(\|x(s_{0})\|_{\infty}+\|x^{*}\|_{\infty})/\theta(0)\leq\sqrt{N}(C_{x^{*}}+C_{x_{0}})/\theta(0)$. If $s_{0}>0$ (i.e., there are DoS attacks since the initial time), by the analysis in Steps 1-3, one can verify that $\bar{y}(s_{0})=\bar{y}(0)$, $x(s_{0})=x(0)$ and $\theta(s_{0})=\theta(0)$. This implies that $\|\pi(s_{0})\|=\|\pi(0)\|$, whose value has been provided in the case of $s_{0}=0$. Overall, for both cases, one has $\|\pi(s_{-1})\|=\|\pi(s_{0})\|\leq\sqrt{N^{2}C_{x_{0}}^{2}+N(C_{x^{*}}+C_{x_{0}})^{2}}/\theta_{0}$.
@@
$D(\delta)$ $:=\frac{dB(\delta)}{d\delta}=-l\sqrt{N}(1-\sqrt{1-2\mu\delta+l^{2}(\delta)^{2}})$
@@
$\quad((1-2\mu\delta+l^{2}(\delta)^{2})(2l^{2}(\delta)^{2}-2\mu)).$ (54)

### v3

@@ V Simulation
For DoS attacks, we consider a sustained DoS attack with variable period and duty cycle, generated randomly. Over a simulation horizon of $1500$s, the DoS signal yields $|\Xi(0,1500)|=1465$s and $n(0,1500)=50$. This corresponds to values (averaged over $1500$s) of $\tau_{D}\approx 30$ and $1/T\approx 0.9767$, and hence $\Delta/\tau_{D}+1/T\approx 0.98$, which approaches the maximum tolerable DoS attack in Theorem 1. In Fig. 1, most of the simulation time is occupied by gray stripes implying the presence of DoS. The white stripes represent DoS-free intervals. Under such a strong level of DoS attacks, one can see in the top plot of Fig. 1 that the actions of the players converge to the NE. In the bottom plot in Fig. 1, one can see that the scaling function $\theta$ converges to zero as time elapses. Moreover, $\theta$ holds its value in the presence of DoS and decreases during DoS-free intervals, i.e., the zooming-in and holding. In Fig. 2, we present the quantized values for $x$ and $y$, i.e., (14) and (13), respectively. One can see that the actual utilized quantization range of $x$ is between $-10$ and $1$, and the one for $y$ is between $-6$ and $6$. They do not encounter saturation problem under DoS attacks and both satisfy the conditions for quantizer unsaturation in (39). One can see that the result in (39b) is conservative. One of the major reasons is the multiplication of $\|G\|$, $\|A_{0}\|$ and $\|S\|$ in (31b) for calculating the value in (39b). Note that constructiveness of quantization levels often occurs in quantized control of multi-agent systems even without DoS attacks (see e.g., [24, 25]). This is majorly due to the nature of distributed systems. If global state is unknown, the agents cannot “well” estimate the states of their neighbors. By contrast, for centralized systems, the problem of constructiveness can be well tackled [22].
@@ VI Conclusions and future research
In the future, it is interesting to explore quantized distributed NE seeking under DoS attacks by Lyapunov-based approaches by following [11]. Extending our results to the players in a general linear form is also a practical direction[23]. It is also useful to apply our results to defend cyber-attacks [2]. At last, we expect that the results in our paper still hold under a directed and strongly connected graph in view of [5].
@@
$\leq\sqrt{1-2\delta\mu+\delta^{2}l^{2}}\|x(k)-x^{*}\|+N\delta l\|\bar{y}(k)\|$ (43)
@@
Therefore, (-B) describes the system dynamics between consecutive successful transmissions at $s_{r}$ and $s_{r+1}$, between which packet losses can occur due to DoS attacks.
@@
In the following, we show that $\|\pi(s_{0})\|$ in (-B) is upper bounded by $\sqrt{N^{2}C_{x_{0}}^{2}+N(C_{x^{*}}+C_{x_{0}})^{2}}/\theta_{0}$. By $\pi(k)=[\|\beta(k)\|\,\,\|\chi(k)\|]^{T}$, it is sufficient to calculate the upper bounds of $\|\beta(s_{0})\|$ and $\|\chi(s_{0})\|$ individually. We present the analysis by the following two cases. If $s_{0}=0$ (i.e., there has no DoS attack at the beginning), we have $\|\beta(s_{0})\|=\|\bar{y}(s_{0})\|/\theta(s_{0})=\|\bar{y}(0)\|/\theta(0)=\|y(0)-\mathbf{1}_{N}\otimes x(0)\|/\theta(0)=\|\mathbf{1}_{N}\otimes x(0)\|/\theta(0)\leq N\|\mathbf{1}_{N}\otimes x(0)\|_{\infty}/\theta(0)=NC_{x_{0}}/\theta(0)$. Similarly, one has $\|\chi(s_{0})\|=\|x(s_{0})-x^{*}\|/\theta(s_{0})\leq(\|x(s_{0})\|+\|x^{*}\|)/\theta(0)\leq\sqrt{N}(\|x(s_{0})\|_{\infty}+\|x^{*}\|_{\infty})/\theta(0)\leq\sqrt{N}(C_{x^{*}}+C_{x_{0}})/\theta(0)$. If $s_{0}>0$ (i.e., there are DoS attacks since the initial time), by the analysis in Steps 1-3, one can verify that $\bar{y}(s_{0})=\bar{y}(0)$, $x(s_{0})=x(0)$ and $\theta(s_{0})=\theta(0)$. This implies that $\|\pi(s_{0})\|=\|\pi(0)\|$, whose value has been provided in the case of $s_{0}=0$. Overall, for both cases, one has $\|\pi(s_{-1})\|=\|\pi(s_{0})\|\leq\sqrt{N^{2}C_{x_{0}}^{2}+N(C_{x^{*}}+C_{x_{0}})^{2}}/\theta_{0}$.
@@
$D(\delta)$ $:=\frac{dB(\delta)}{d\delta}=-l\sqrt{N}(1-\sqrt{1-2\mu\delta+l^{2}\delta^{2}})$
@@
$\quad((1-2\mu\delta+l^{2}\delta^{2})(2l^{2}(\delta^{2}-2\mu)).$ (54)